# AIDE: Task-Specific Fine Tuning with Attribute Guided MultI-Hop Data Expansion

### v1

@@ 4.1 Experiment Setup
Datasets. We evaluate all methods on $5$ tasks from BIG-Bench, $5$ tasks from MMLU, ARC-Challenge, GSM8K, and TruthfulQA. We show the data statistics about the selected tasks in Table 4. Details of the benchmarks are provided in Appendix A.
Base Model Fine-tuning with Data Source BIG-Bench Avg.
@@
Base Model 46.7% 47.7% 61.1% 11.6% 1.4% 33.7%
Llama-3.1-8B AIDE (Ours) 85.0% 98.7% 51.6% 27.7% 87.1% 70.0%
Base Model 63.3% 74.5% 47.8% 21.3% 21.4% 45.6%
Llama-3.2-3B AIDE (Ours) 81.7% 86.3% 50.0% 24.2% 40.8% 56.6%
Base Model 38.3% 32.7% 42.0% 20.2% 17.3% 30.1%
Table 3: We compare the performance of different zero-shot learning methods across $5$ tasks from BIG-Bench. The table setup is similar to that of Table 2.
Baselines. We use pretrained Mistral-7B, Llama-3.1-8B, and Llama-3.2-3B as baselines for comparison with the models fine-tuned using synthetic data from AIDE. We also compare AIDE with SOTA synthesis methods (Evol-Instruct, DataTune, and Prompt2Model) by fine-tuning Mistral-7B. When human-generated (gold) data is available, it serves as another baseline. A fine-tuned Mistral-7B using 250K synthetic data from Evol-Instruct  [^1] https:/huggingface.co/dreamgen/WizardLM-2-7B is utilized as Mistral-7B with Evol-Instruct.
Metrics. We evaluated all models using zero-shot accuracy as the primary metric on the benchmarks. For GSM8K, we report the average 8-shot performance using maj@8, using prompts from Wang et al. (2023). Avg. $\Delta(\uparrow)$ denotes average improvements of models compared to the base models.
Data Synthesis Setup. We configure SOTA data synthesis methods using their default settings. Since BIG-Bench lacks a training set, we sample $10$ task-related seed data points per task from Hugging Face datasets to generate synthetic data. Similarly, for other benchmarks, we sample $10$ seed data points per task from their respective training sets to produce synthetic data. In AIDE, we set the hop depth to $K=2$ and employ Claude Sonnet 3.5 as the LLM extractor, synthesizer, grader and annotator. The LLM extracts ${\mathcal{A}}^{(K)}$ for each data point $X^{(K)}_{i}$, comprising $3$ triplets, each consisting of $1$ topic and $1$ relevant attributes. Each triplet in ${\mathcal{A}}^{(K)}$ is followed by $3$ operations: concretizing, adding constraint and adding reasoning. To further diversify attributes, we retrieve the top-$5$ related personas for each topic. In Table 4, we demonstrate the amount of seed data used and the quantity of data synthesized in AIDE. Specifically, using $K=2$ and $10$ seed data points for each task, AIDE generates approximately $3$K new data points in about $20$ hours when adapting the self-reflection mechanism to improve the quality of new data.
Benchmarks Task Name Abbr. # Test Data # Synthetic Data
BIG-Bench Code Line Descriptions Code 60 3.0K
Cause and Effect C&E 153 3.2K
Implicatures Impl. 492 3.1K
Elementary Math Math 7,688 3.1K
Temporal Sequence Time 1,000 3.2K
MMLU High School Biology Bio. 300 3.4K
College Computer Science CS 100 3.2K
Philosophy Phi. 311 3.4K
Electrical Engineering EE 145 3.0K
Marketing Market. 234 3.3K
ARC-Challenge - - 1,170 3.3K
GSM8K - - 1,320 3.2K
TruthfulQA - - 817 3.1K
Table 4: Statistics of tasks and synthetic data related to BIG-Bench, MMLU, ARC-Challenge, GSM8K and Truthful QA. Note that we employ a self-reflection mechanism to improve the quality of synthetic data, which filters out lower-quality synthetic data. Using $K=2$ and $10$ seed data points per task, AIDE synthesizes approximately $3$K data points for each task.
Fine-tuning Setup. We applied the LoRA Hu et al. (2022) to fine-tune Mistral-7B. We randomly split $10\%$ of the synthetic data as validation set while the rest of synthetic data as training set. The process was carried out over $10$ epochs with batch size equal to $10$. We select learning rate $55e-$ 5 with LoRA’s $\alpha$ parameter as $16$ and choose the run with the lowest validation loss at any point. We used the AdamW optimizer Loshchilov and Hutter (2019) and set LoRA $r=8$. We conduct our training on a server with $8$ NVIDIA A100 GPUs.
@@ 4.2 Performance and Analysis (RQ1)
Table 5: Different core components of AIDE contribute to the synthetic data, improving the performance of Mistral-7B on the Time task, one of the most challenging task in the BIG-Bench. We highlight the best performance and the base performance is in Table 3.
In Table 2, on MMLU tasks, Mistral-7B fine-tuned with AIDE data achieves $18.9\%$ improvement over the base model, outperforming Mistral-7B fine-tuned with gold training data and Evol-Instruct by $1.7\%$ and $13.8\%$, respectively. On TruthfulQA, AIDE achieves the best performance compared to the base Mistral-7B and other fine-tuned Mistral-7B variants. On ARC-Challenge, Mistral-7B fine-tuned with AIDE is the runner-up, surpassing base models and improving over Evol-Instruct by $>20.0\%$. On GSM8K, the performance is comparable to Evol-Instruct. Fine-tuning on AIDE also improves the average accuracy of Llama-3.1-8B by $10.8\%$ and Llama-3.2-3B by $10.8\%$.
In Table 3 (BIG-Bench without training sets), Mistral-7B with AIDE significantly outperforms itself fine-tuned using Prompt2Model and DataTune by $>38.0\%$, and its base model by $>40.0\%$. Similarly, fine-tuning Llama-3.1-8B and Llama-3.2-3B with synthetic data from AIDE results in $>20.0\%$ average accuracy improvement.
4.3 Ablation and Sensitivity Studies (RQ2)
We conduct ablation studies to empirically explore AIDE with Mistral-7B as the base model.
Figure 4: The effect of different number of seed data w/ and w/o task demonstration on the Time task from BIG-Bench.
@@ 4.3 Ablation and Sensitivity Studies (RQ2)
Figure 5: Effect of data quantity with different number of $K$ values in multi-hop synthesis. The different tasks are from the BIG-Bench benchmark.
@@
Scaling with Data Quantity using Different Depth $K$. The multi-hop depth $K$ determines the amount of AIDE’s synthetic data, directly influencing fine-tuned model performance. Figure 5 shows increasing $K$ from 2 to 4 significantly enhances Mistral-7B’s performance on the code task after fine-tuning on AIDE data. However, for other tasks, performance gains gradually decrease with higher $K$ values due to the inherent ability gap between the base model and the LLM synthesizer.
@@
Table 6: Comparison of the performance of Mistral-7B fine-tuned with synthetic data from AIDE. We highlight the best performance.
Effect of Capability of LLMs. We investigate the impact of using different LLMs as components in AIDE by conducting experiments on $5$ BIG-Bench tasks, using Claude Sonnet 3.5 and GPT-3.5-Turbo separately to synthesize data. As shown in Table 6, fine-tuning Mistral-7B with AIDE’s synthetic data, generated with either Claude Sonnet 3.5 or GPT-3.5-Turbo as components, enhances the base model Mistral-7B’s performance by $>40.0\%$.
4.4 Analysis of Relevance, Diversity and Complexity (RQ3)
We empirically investigate the relevance, diversity and complexity of synthetic data from AIDE.
4.4.1 Analysis of Relevance
Figure 6: The relevance score related to the sampled synthetic data and task-related seed data from the Code, C&E and Impl. tasks.
(a) Synthetic data (Code task)
(b) Synthetic data (C&E task)
(c) Synthetic data (Impl. task)
@@ 4.4.1 Analysis of Relevance
Since seed data is task-relevant, synthetic data should also exhibit task relevance if it closely aligns with the seed data. To evaluate this, we randomly sampled $10$ synthetic data points per task from the Code, C&E, and Impl. tasks in BIG-Bench. Using the jina embedding model Günther et al. (2023) to encode all data points, we calculate the similarity between the embeddings of the synthetic data and their corresponding seed data. The results, depicted in Figure 7(a), show that the similarity have an average exceeding $0.5$, indicating synthetic data has strong relevance to the seed data. Similarly, Figures 7(b) and 7(c) reveal that synthetic data from the C&E and Impl. tasks exhibit even higher average similarity scores, surpassing $0.6$. These findings demonstrate that the synthetic data is highly relevant to the respective tasks.
Additionally, using the prompt shown in Figure 19, we employ Claude Sonnet 3.5 to evaluate the relevance scores between synthetic data and their corresponding seed data. Claude Sonnet 3.5 assigns scores on a scale from 0 to 10, with 10 indicating the highest relevance. As illustrated in Figure 6, the average relevance scores across the three tasks, along with their standard deviations, predominantly range from 5 to 9. These results further confirm the strong task relevance of the synthetic data.
4.4.2 Analysis of Diversity
@@
(a) Synthetic data (Code task)
(b) Synthetic data (C&E task)
(c) Synthetic data (Impl. task)
@@
Figure 8: The diversity of knowledge from randomly sampling $20$ synthetic data synthesized by AIDE for the Code task, the C&E task and the Impl. task from the BIG-Bench.
@@
To assess the diversity of the synthetic data, we randomly sample $20$ synthetic data points per task from the Code, C&E, and Impl. tasks. Using the prompt shown in Figure 18, we apply Claude Sonnet 3.5 to analyze the diversity of the synthetic data based on related knowledge, as depicted in Figure 8. In Figure 8(a), the sampled synthetic data related to coding includes various programming topics and operations. For the C&E and Impl., the synthetic data references a wide range of different knowledge areas, as shown in Figures 8(b) and 8(c).
Furthermore, following the approach in Ye et al. (2022a), we compute Self-BLEU Zhu et al. (2018) to quantitatively evaluate the diversity of both synthetic and gold data. For tasks without training sets, we use their test sets as the gold data. Self-BLEU is calculated by computing the BLEU score of each synthetic data using all other synthesis data in the evaluation set as references. A lower Self-BLEU score indicates higher diversity. Specifically, we randomly sample $300$ data from synthetic data and gold data respectively, comparing each data point to all others. In tasks with fewer than $300$ gold data points, we sample $10\%$ of the gold data from the tasks to compare with all other data during the computation of Self-BLEU. We report 4-gram based Self-BLEU in Table 7. The results reveal that synthetic data from AIDE achieves Self-BLEU scores comparable to those of gold data for most tasks, demonstrating AIDE’s ability to generate diverse synthetic data.
4.4.3 Analysis of Complexity
(a) Synthetic data (Code task)
(b) Synthetic data (C&E task)
(c) Synthetic data (Impl. task)
@@ 4.4.3 Analysis of Complexity
Similar to Evol-Instruct (Xu et al., 2024) using $5$ predefined operations to expand the complexity of synthetic data, AIDE utilizes $3$ predefined operations including reasoning, constraint and concrete following triplets from ${\mathcal{A}}$ to expand the complexity during data synthesis. To evaluate the complexity of synthetic data generated by AIDE, we randomly sample $500$ data points from various synthesis methods, including Alpaca, Evol-Instruct, and AIDE. We then use Claude Sonnet 3.5 to evaluate the complexity of the synthetic data, applying the same prompt used in Evol-Instruct. We plot the distribution of score of complexity ranging from $1$ to $10$, shown on Figure 9. We find that most of synthetic data from AIDE and Evol-Instruct achieve complexity scores above $5$, in contrast to the data from Alpaca. It is worth mentioning that AIDE only uses $3$ predefined operations, fewer than five operations employed in Evol-Instruct, yet still produces synthetic data of comparable complexity.
4.4.4 Visualization
@@
Following the approach in Zhao et al. (2024b), we analyze the coverage of synthetic data from AIDE in the embedding space. Specifically, we use the jina-embeddings-v2-base-codeGünther et al. (2023) to embed data points related to coding, while employing jina-embeddings-v2-base-en to encode other text data. With these embeddings, we utilize t-SNE van der Maaten and Hinton (2008) to project embeddings into a two-dimensional space. For comparison, we adopt the real data from the code and C&E tasks as baselines to demonstrate the coverage of synthetic data from AIDE.
In Figure 10(a), we observe that the embedding clusters of synthetic data generated by AIDE for the Code task largely overlap with the embeddings of all real data from the same task. Figure 10(b) shows that the synthetic data spans a broader range, covering all real data from the C&E task. This suggests that AIDE, even with a small number of seed data points related to specific tasks, effectively captures diverse distributions within the target task space. As a result, fine-tuning Mistral-7B with synthetic data from AIDE has a positive impact on the model’s performance across specific tasks.
5 Conclusion
Existing data synthesis methods struggle to produce synthetic data that is both task-relevant and diverse for fine-tuning, or require large seed datasets. In this paper, we introduce AIDE, a novel framework that enables task-relevant, diverse, and high-quality data expansion from few seed examples. It features multi-hop synthesis guided by attributes and personas, a residual connection to reduce irrelevance at deeper hops, and self-reflection to enhance quality. Our experimental results show that fine-tuning Mistral-7B and Llama models on AIDE’s synthetic data significantly improves their performance over base models and surpasses SOTA synthesis methods. AIDE addresses key limitations in generating effective task-specific training data.
Appendix A Benchmark Statistics

### v2

@@ 4.1 Experiment Setup
Datasets. We evaluate all methods across 5 tasks from BIG-Bench, 5 tasks from MMLU, 1 task from FinBen, as well as MedQA, ARC-Challenge, GSM8K, and TruthfulQA. Details of the benchmarks and statistics of the synthetic data from AIDE are provided in Appendix H and F.
Pretrained Model Fine-tuning with Data Source BIG-Bench Avg. ($\uparrow$)
@@
Pretrained Model 46.7% 47.7% 61.1% 11.6% 1.4% 33.7%
Table 3: AIDE vs. SOTA Data Synthesis Methods. We compare the performance of various zero-shot learning approaches in Mistral-7B fine-tuned with AIDE and SOTA synthesis methods across five BIG-Bench tasks. The table follows a setup similar to Table 2. Notably, Evol-Instruct fine-tunes Mistral-7B with 250K synthetic data points.
Baselines. We use fine-tuned Mistral-7B, Llama-3.1-8B, and Llama-3.2-3B with human-generated (gold) data as baselines for comparison with the models fine-tuned using synthetic data from AIDE. We also compare AIDE with SOTA synthesis methods (Evol-Instruct, DataTune, and Prompt2Model) by fine-tuning Mistral-7B. A fine-tuned Mistral-7B using 250K synthetic data from Evol-Instruct  [^1] https:/huggingface.co/dreamgen/WizardLM-2-7B is utilized as Mistral-7B with Evol-Instruct. Details about the setups are provided in Appendix E.
Metrics. We evaluated all models using zero-shot accuracy as the primary metric on the benchmarks. For GSM8K, we report 8-shot maj@8 performance using prompts from Wang et al. (2023).
@@ 4.2 Performance and Analysis (RQ1)
Table 4: Different core components of AIDE contribute to the synthetic data, improving the performance of Mistral-7B on the Time task from BIG-Bench. We highlight the best performance and the base performance is in Table 3.
In Table 2, the pretrained models fine-tuned with AIDE demonstrate comparable or superior performance to those fine-tuned with gold data. For example, on MMLU tasks, models fine-tuned with AIDE data outperform those trained on gold data by an average of $>1.4\%$. In the CFA task, synthetic data from AIDE improves Mistral-7B and Llama-3.1-8B by at least $>1.6\%$ compared to gold data. On ARC-Challenge, the Llama series fine-tuned with AIDE surpasses their counterparts fine-tuned on gold data. In GSM8K, pretrained models fine-tuned with AIDE perform comparably to those fine-tuned with gold data. On TruthfulQA, models fine-tuned with AIDE exceed those trained on gold data by an average of $>15.0\%$. Similarly, on MedQA, AIDE improves pretrained models by more than $>8.2\%$ on average. In Table 3 (BIG-Bench without training sets), Mistral-7B with AIDE significantly outperforms itself fine-tuned using Evol-Instruct, Prompt2Model and DataTune by $>20.0\%$, and its pretrained model by $>40.0\%$. This is because Prompt2Model focuses on generating task-specific data with limited diversity, whereas Evol-Instruct, despite its multi-hop synthesis structure, generates data without targeting a specific task.
Figure 2: The effect of varying the number of seed data w/ and w/o task demonstration on the Time task from BIG-Bench.
@@
Figure 3: Effect of data quantity with different number of $K$ values in multi-hop synthesis based on the BIG-Bench.
4.3 Ablation and Sensitivity Studies (RQ2)
We conduct ablation studies to empirically explore AIDE with pretrained models.
@@ 4.3 Ablation and Sensitivity Studies (RQ2)
Scaling with Data Quantity using Different Depth $K$. The multi-hop depth $K$ determines the amount of AIDE’s synthetic data, directly influencing fine-tuned model performance. Figure 3 shows increasing $K$ from 2 to 4 significantly enhances Mistral-7B’s performance on the code task after fine-tuning on AIDE data. However, for other tasks, performance gains gradually decrease with higher $K$ values due to the inherent ability gap between the pretrained model and the LLM synthesizer.
Effect of Residual Connection. We use a contract task from LegalBench Guha et al. (2023), setting the hop depth $K$ to 4 while varying the depth of residual connections $L$. By synthesizing 5,682 training data points from 6 seeds, we analyze their impact on fine-tuning models. Figure 4 shows that as the multi-hop synthesis depth increases, a higher residual connection depth $L$ improves the task relevance of the synthetic data, resulting in better model performance during fine-tuning.
Figure 4: The effect of varying the depth of residual connections ($L$) when we fix the hop depth $K$ as 4.
Effect of Capability of LLMs. We investigate the impact of using different LLMs as components in AIDE by conducting experiments on $5$ BIG-Bench tasks, using Claude Sonnet 3.5 and GPT-3.5-Turbo separately to synthesize data. As shown in Table 5, fine-tuning Mistral-7B with AIDE’s synthetic data, generated with either Claude Sonnet 3.5 or GPT-3.5-Turbo as components, enhances the pretrained model Mistral-7B’s performance by $>40.0\%$.
@@
Table 5: The performance of Mistral-7B fine-tuned with synthetic data from AIDE using different LLMs as synthesizer.
4.4 Relevance and Diversity (RQ3)
We empirically investigate the relevance and diversity of synthetic data from AIDE. Appendix G provides details on synthetic data complexity.
Figure 5: The relevance score related to the sampled synthetic data and task-related seed data from the Code task, the C&E task and the Impl. task.
(a) Synthetic data for the Code task
(b) Synthetic data for the C&E task
(c) Synthetic data for the Impl. task
@@ 4.4 Relevance and Diversity (RQ3)
(a) Synthetic data for the Code task
(b) Synthetic data for the C&E task
(c) Synthetic data for the Impl. task
@@
Figure 7: We assess the diversity of knowledge by randomly sampling 20 synthetic data points generated by AIDE for the Code, C&E, and Impl. tasks from BIG-Bench.
Analysis of Relevance. Since the seed data is task-specific, the synthetic data should also be task-relevant if it closely aligns with the seed data. To evaluate this, we randomly sample 10 synthetic data points per task from the Code, C&E, and Impl. tasks in the BIG-Bench benchmark. We use the Jina embedding model Günther et al. (2023) to encode all data points, and compute the similarity between each synthetic data point and its corresponding seed data. As shown in Figure 6, the synthetic data exhibits strong relevance to the seed data, with an average similarity score above 0.5.
Additionally, we employ Claude Sonnet 3.5 to assess the relevance of synthetic data to the seed data across the three tasks. Claude assigns a relevance score from 0 to 10, with 10 indicating the highest relevance. As shown in Figure 5, the average scores range from 5 to 9, further confirming the task alignment of the synthetic data. The standard deviation arises because the samples contain data points with significant diversity, yet remain relevant to the corresponding task.
@@
Analysis of Diversity. AIDE expands attributes through using topics to retrieve personas from Persona Hub, which diversifies the data synthesis. To verify the diversity of synthetic data, we randomly sample $20$ synthetic data per task from the Code, C&E, and Impl. tasks. Using the prompt shown in Figure19, we employ Claude Sonnet 3.5 to assess the diversity of the synthetic data based on relevant knowledge. As illustrated in Figure 7(a), the sampled synthetic data for the Code task covers a variety of programming topics and operations. In the C&E and Impl. tasks, we observe that the synthetic data spans a wide range of knowledge domains, as shown in Figures 7(b) and 7(c).
Additionally, following prior work Ye et al. (2022a), we compute Self-BLEU Zhu et al. (2018) to quantitatively assess the diversity of both synthetic and gold data. The results in Table 6 show that the synthetic data generated by AIDE achieves Self-BLEU scores comparable to those of gold data across most tasks, demonstrating its effectiveness in producing diverse synthetic data.
5 Conclusion
Existing data synthesis methods struggle to generate synthetic data that is both task-relevant and diverse for fine-tuning or require large seed datasets. In this paper, we introduce AIDE, a novel framework that enables task-relevant, diverse, and high-quality data expansion from few seed examples. It features multi-hop synthesis guided by attributes and personas, along with a residual connection to mitigate irrelevance at deeper hops. Our experiments show that fine-tuning Mistral-7B and Llama models with AIDE outperforms the models fine-tuned with gold data and SOTA synthesis methods.
Appendix A Detailed Related Work
Appendix B Multi-Hop Synthesis
The Figure 8 shows an example of the multi-hop synthesis, which the seed data $X^{(0)}_{i}$ is used to synthesize its $1$-hop neighbors $X^{(1)}_{1}$ and $X^{(1)}_{2}$ during the $1$-hop synthesis. Similarly, each $1$-hop neighbor can be applied to generate $2$-hop neighbors of $X^{(0)}_{i}$. For each input data $X^{(0)}_{i}$ where $1\leq i\leq n$, we recursively synthesis data using the same pattern until reaching the depth of $K$.
Appendix C An Example of Unfolded Multi-Hop Synthesis
Figure 9: An example of unfolded multi-hop synthesis when $K=2$.
Figure 9 illustrates an example of unfolded multi-hop synthesis. In this example, we set $K=2$. $X^{(0)}_{i}$ is one of the seed data point and $X^{(1)}=\{X^{(1)}_{1},X^{(1)}_{2},...,X^{(1)}_{m_{1}}\}$ represents synthetic data from $1$-hop synthesis while $X^{(2)}=\{X^{(2)}_{1},X^{(2)}_{2},...,X^{(2)}_{m_{2}}\}$ represents synthetic data from $2$-hop synthesis. $r$ is the relation between a topic $t$ and knowledge attribute $a$. The predefined operation $Op$ is the abbreviation of operation. Green area includes a path of synthesis showing the relevance between two data points. Orange area shows a path to synthesize data with diversity and relevance. We zoom in one of the branches related to $X^{(1)}_{3}$ in $2$-hop synthesis. Table 1 demonstrates an example of the synthesis.
Appendix D Residual Connection
We introduce residual connections between a seed data point and its neighbors. Specifically, for any depth $d$ where $1<d\leq K$, we establish connections when $d\leq L$ where $L$ is the depth of residual connection within the range $(1,K]$. For example, in Figure 9, when $K=2$, setting $L=2$ allows connections between the seed data and all neighbors at hop depth 2, ensuring seed information is available for generating the neighbors.
Experiments in Figure 4 demonstrate that when the hop depth $K$ is large, applying residual connections with a greater depth $L$ enhances the relevance of the synthetic data, leading to improved performance in the fine-tuned model. However, as hop depth $K$ increases, removing low-relevance neighbors instead of using residual connections to retain them can lead to a reduction in the amount of synthetic data.
Appendix E Detailed Experimental Setup
Data Synthesis Setup. We configure the SOTA data synthesis methods using their default settings. Since BIG-Bench lacks a training set, we sample $10$ task-related seed data points per task from Hugging Face datasets to generate synthetic data. For the remaining benchmarks, we similarly sample $10$ seed data points per task from their respective training sets to produce synthetic data. We set the depth of hop $K=2$ in the multi-hop synthesis. We employ Claude Sonnet 3.5 as the LLM generator, the LLM synthesizer, the LLM grader and the LLM annotator in AIDE. We require the LLM to generate ${\mathcal{A}}^{(K)}$ of a data point $X^{(K)}_{i}$, which consists of $1$ topic and $3$ most related attributes. Each triplet in ${\mathcal{A}}^{(K)}$ followed by $3$ operations: concretizing, adding constraint and adding reasoning. With a topic, we retrieve top-$5$ related personas to diversify attributes.
Fine-tuning Setup. We applied the LoRA (Hu et al., 2022) to fine-tune Mistral-7B. We randomly split $10\%$ of the synthetic data as validation set while the rest of synthetic data as training set. The process was carried out over $10$ epochs with batch size equal to $10$. We select learning rate $55e-$ 5 with LoRA’s $\alpha$ parameter as $16$ and choose the run with the lowest validation loss at any point. We used the AdamW optimizer (Loshchilov and Hutter, 2019) and set LoRA $r=8$. We conduct our training on a server with $8$ NVIDIA A100 GPUs.
Self-Reflection for Synthetic Data To ensure the correctness, relevance, and diversity of synthetic data, we apply existing self-reflection techniques (Madaan et al., 2023; Pan et al., 2024) after synthesis (Figure 1). A LLM grades synthetic data $X^{(K)}_{i}$ on these aspects, providing a score (from $1$ to $10$) and feedback. Data exceeding a score threshold (i.e., threshold equal to $5$) is added to the dataset; otherwise, it undergoes limited re-synthesis iterations. A LLM annotator then labels the data, with self-reflection ensuring labeling correctness. Related prompts are shown in Appendix N.
Appendix F Statistics of Synthetic Data
Benchmarks Task Name Depth of $K$ Amount of seed Data Quantity of Synthetic Data
BIG-Bench Code 2 10 3.0K
C&E 2 10 3.2K
Impl. 2 10 3.1K
Math 2 10 3.1K
Time 2 10 3.2K
MMLU Bio. 2 10 3.4K
CS 2 10 3.2K
Phi. 2 10 3.4K
EE 2 10 3.0K
Market. 2 10 3.3K
ARC-Challenge - 2 10 3.3K
GSM8K - 2 10 3.2K
TruthfulQA - 2 10 3.1K
FinBen CFA 2 10 893
MedQA - 2 10 2.2K
Table 7: Statistics of synthetic data. Note that we adapt the self-reflection mechanism to enhance data quality, which also filters out some synthetic data.
In Table 7, we demonstrate the amount of seed data used and the quantity of data synthesized in AIDE. Specifically, using $K=2$ and $10$ seed data points for each task, AIDE generates approximately $3$K new data points in about $20$ hours when adapting the self-reflection mechanism to improve the quality of new data.
Appendix G Detailed Analysis of Relevance, Diversity and Complexity (RQ3)
We conduct experiments to assess whether the synthetic data generated by AIDE preserves its complexity.
G.1 Analysis of Complexity
(a) Synthetic data for the Code task
(b) Synthetic data for the C&E task
(c) Synthetic data for the Impl. task
@@ G.1 Analysis of Complexity
Similar to Evol-Instruct (Xu et al., 2024) using $5$ predefined operations to expand the complexity of synthetic data, AIDE utilizes $3$ predefined operations including reasoning, constraint and concrete following triplets from ${\mathcal{A}}$ to expand the complexity during data synthesis. For verifying the complexity of synthetic data from AIDE, we randomly sample $500$ synthetic data from different synthetic methods including Alpaca, Evol-Instructs and our AIDE. Then we apply Claude Sonnet 3.5 to evaluate the complexity of synthetic data using the same prompt as that from Evol-Instruct. We plot the distribution of score of complexity ranging from $1$ to $10$, shown on Figure 10. We find that most of synthetic data from AIDE and Evol-Instruct obtain the score of complexity higher than $5$, when comparing with that from Alpaca. It is worth mentioning that AIDE only uses $3$ predefined operations less than the operations applied in Evol-Instruct while having the synthetic data with comparable complexity.
@@
G.2 Visualization
We follow the approach in Zhao et al. (2024b) and analyze the coverage of synthetic data from AIDE in the embedding space. Specifically, we use the jina-embeddings-v2-base-code Günther et al. (2023) to embed data points about coding while employ jina-embeddings-v2-base-en to encode other text data. With the embeddings, we utilize t-SNE van der Maaten and Hinton (2008) to project embeddings into a two-dimensional space. We adopt the real data from the code line description task and the C&E task as baselines to demonstrate the coverage of synthetic data from AIDE.
In Figure 11(a), we observe that the embedding clusters of synthetic data via AIDE and the embeddings of all real data from the Code task appear to be largely disjoint. Figure 11(b) demonstrates that the synthetic data has a larger range which covers all real data from the C&E task. This supports a conclusion that AIDE with few seed data related to specific tasks systematically cover different distributions of the target task space, and therefore fine-tuning Mistral-7B with synthetic data from AIDE leads to a positive effect on the improvement of performance of Mistral-7B in specific tasks.
Appendix H Benchmark Statistics